# TechCoach: Towards Technical Keypoint-Aware Descriptive Action Coaching

### v1

@@ Abstract
To guide a learner to master the action skills, it is crucial for a coach to 1) reason through the learner’s action execution and technical keypoints, and 2) provide detailed, understandable feedback on what is done well and what can be improved. However, existing score-based action assessment methods are still far from this practical scenario. To bridge this gap, we investigate a new task termed Descriptive Action Coaching (DAC) which requires a model to provide detailed commentary on what is done well and what can be improved beyond a quality score from an action execution. To this end, we construct a new dataset named EE4D-DAC. With an LLM-based annotation pipeline, our dataset goes beyond the existing action assessment datasets by providing the hierarchical coaching commentary at both keypoint and instance levels. Furthermore, we propose TechCoach, a new framework that explicitly incorporates keypoint-level reasoning into the DAC process. The central to our method lies in the Context-aware Keypoint Reasoner, which enables TechCoach to learn keypoint-related quality representations by querying visual context under the supervision of keypoint-level coaching commentary. Prompted by the visual context and the keypoint-related quality representations, a unified Keypoint-aware Action Assessor is then employed to provide the overall coaching commentary together with the quality score. Combining all of these, we build a new benchmark for DAC and evaluate the effectiveness of our method through extensive experiments. Data and code will be publicly available.
footnotetext: *: Corresponding author. $\dagger$: Equal contributions.
@@ 1 Introduction
Figure 1: An illustration of real-world action coaching. Upper: Given an action demonstration, a human coach will provide coaching comments on what is done well and what can be improved. Lower: To provide precise feedback, the coach incorporates Keypoint-level Reasoning into the action coaching process, i.e., keep the technical keypoints (e.g., The player should extend the arms fully as they bring the ball under the basket) in mind and determine the keypoint-level action quality (e.g., The athlete performs beautifully with the right arm extension) of the action. Such a reasoning process motivates us to propose our TechCoach to address Descriptive Action Coaching.
@@
- Functionality. In practical action-guiding scenarios, a coach must provide detailed and understandable feedback on what is done well and what can be improved so that the learner can be fully aware of the execution details and master the skills (see the upper part of Fig. 1). However, existing works pose AQA as a score-regression [28, 35, 50] or pairwise-ranking [9, 55, 12] problem, limiting the applicability of the models to more practical scenarios.
@@
- Reasoning process. To provide precise feedback, a coach will keep the technical keypoints (e.g., The player should extend the arms fully in Reverse Layup) in mind and determine the keypoint-level action quality from action execution (see the lower part of Fig. 1). However, few existing works have incorporated such a keypoint-level reasoning process into AQA. While recent works [39, 15] propose to address this issue by recognizing the existence of keypoint-level mistakes, such a paradigm still limits deeper exploration of the connections between the action execution and the keypoint-level action quality (i.e., detailed strength/weakness on a technical keypoint).
@@
To bridge the gap between the field of AQA and real-world action coaching scenarios, we investigate a new task termed Descriptive Action Coaching (DAC), which requires the model to provide detailed commentary on what is done well and what can be improved, beyond merely assigning a quality score to an action execution. Such a task is challenging because to provide such a detailed commentary, the model not only needs to understand fine-grained action execution (e.g., the player looks down when performing Reverse Layup), but also build explicit connections between the execution with the action quality (e.g., it is a weakness to look down at the ball as it will miss the target).
@@
To facilitate the studies on DAC, we construct EE4D-DAC, a new dataset that not only contains various action videos and quality scores but also features hierarchical detailed coaching commentary on both keypoint and instance levels. Specifically, we first source the action videos from the recently proposed EgoExo4D [10] dataset. Subsequently, an LLM-driven automatic annotation pipeline is proposed to progressively obtain the general technical keypoints and hierarchical coaching commentary.
@@
The construction of the EE4D-DAC dataset allows us to build Technical Keypoint-aware Descriptive Action Coach (TechCoach), a new DAC framework which explicitly incorporates keypoint-level reasoning into the action coaching process. The key idea of TechCoach is to imitate the real-world coaching scenario, where the coach first reasons the action quality on each technical keypoint, and then provides feedback by incorporating the keypoint-level reasoning results and visual context.
@@
To this end, TechCoach includes a Context-aware Keypoint Reasoner, which queries the visual context with general technical keypoints to obtain keypoint-related quality representations. To ensure such representations carry the action quality information, a keypoint-level alignment loss is proposed to align the representations with the keypoint-level coaching commentary. After that, a Unified Keypoint-aware Action Assessor (KA2) is utilized to regress the action quality score and generate overall coaching commentary. In this process, a progressive action coaching attention mask is adopted to guide KA2 to integrate information starting with the visual context level, followed by the keypoint level, and ultimately the decision (i.e., score-prediction and commentary-generation) level.
@@
Combining the EE4D-DAC dataset and TechCoach, we can construct a new benchmark to expand the field of AQA. Extensive experiments demonstrate that: 1) Our TechCoach achieves state-of-the-art performance among the compared methods, including task-specific models and general MLLMs; 2) The Context-aware Keypoint Reasoner is a core design and cannot be simply replaced by other solutions (e.g., replacing the Keypoint-level Alignment with a classification-based solution that classify whether the execution shows strength or weakness on each keypoint).
@@
In summary, the contributions of this work can be summarized as: 1) We construct a new dataset termed EE4D-DAC to address Descriptive Action Coaching. To our knowledge, EE4D-DAC is the first dataset that includes hierarchical coaching commentary on both keypoint level and instance level. 2) We develop TechCoach, a new method to address DAC by explicitly collaborating keypoint-level reasoning into the action coaching process. 3) We construct a new benchmark on DAC. Experiments not only evaluate the effectiveness of the proposed approach but also provide a starting point for future work in this field.
@@
Figure 2: An overview of the LLM-driven automatic annotation pipeline for the EE4D-DAC dataset. We first prompt the LLM to generate general technical keypoints for the given action task (Orange $\rightarrow$). Subsequently, we ask the LLM to summarize the keypoint-level and instance-level commentary according to the general technical keypoints and the original expert commentary (Blue $\rightarrow$). Information about various keypoints is highlighted in colors. Zoom in for best view.
@@ 2 Related Works
Action Quality Assessment (AQA) [43, 35, 50, 24, 47, 45, 3, 44, 29, 51, 27, 54, 30, 15, 22, 21, 57] aims at evaluating how well an action is performed. Most existing works pose AQA as a score regression task, and approach this task in many ways, e.g., introducing new formulations [43, 35, 50], augmenting human-body representations [24, 47], exploring temporal structures[45, 3, 44], collaborating multi-modal information [29, 51, 42]. Although impressive progress has been achieved, regressing a single score is still far from real-world action coaching scenarios. While [27, 54] propose to generate commentary in addition to the score, the commentary comes from the live comment, focusing on action and score rather than skill guiding and coaching. Recently, considering it is crucial for a coaching model to reason through various technical keypoints, [30, 15, 39] propose to detect mistakes on various body parts or technical keypoints. [22, 21] introduce a scoring rubric to inject judging rules into the score prediction process. However, This branch of methods is still limited by the binary or score-only feedback. Notably, [23] incorporates multiple models together with judging rules to provide formatted feedback for diving judging, but requires complex manual design on a single task. Differently, our work focuses on generating free-from commentary on what is done well and what can be improved, which is more aligned with practical coaching scenarios. To achieve this, we construct a new dataset that features detailed coaching commentary on both keypoint and instance levels. Furthermore, we propose a new general framework that collaborates key-point reasoning into coaching process.
@@
Video Captioning [34, 46, 17, 40, 19, 49] is a long-lasted task that requires a model to generate language caption with video frames as input. Early studies focus on designing stronger architectures [17, 40, 1, 33, 48] or utilizing video-language pretraining [34, 19, 49, 2, 32] to model relationships between video and language. Recently, Multimodal Large Language Models (MLLMs) [16, 41, 14, 20, 13, 7] show strong generalization ability across various video-language tasks such as video content description and video question answering. In this paper, we focus on Descriptive Action Coaching, a rarely explored but challenging scenario that requires video-language models to collaborate fine-grained video content and domain knowledge to provide detailed coaching commentary. We provide a new testbed for today’s video-language models and show that they still struggle to handle such a practical scenario. Besides, we propose a new keypoint-guided coaching model and evaluate its effectiveness through extensive experiments.
3 EE4D-DAC Dataset
@@
To enable studies on Descriptive Action Coaching, we construct a new dataset termed EE4D-DAC. Different from existing action assessment datasets, EE4D-DAC not only contains various action videos and quality scores but also features hierarchical detailed coaching commentary on both keypoint level and instance level. We will introduce the data source in Sec. 3.1, the annotation pipeline in Sec. 3.2, and provide comparisons with existing datasets in Sec. 3.3 respectively.
@@ 3.1 Data Source
We source data from the recently proposed EgoExo4D [10] dataset. Except for various ego-exo action videos, EgoExo4D also contains execution rating (from 1 to 10) and time-aligned free-form spoken commentary (e.g. “Yeah, see how she’s having to look down. You can’t see the basket here.") by domain experts. Such characteristics make the EgoExo4D dataset a great starting point for studying action coaching.
@@
Datasets Years Data Unique Unique KeyPoint Level Judgment Instance Level Judgment
@@
MIT-Skate [31] ECCV14 Olympics 150 7.3h $\times$ - - $\times$ -
AQA-7 [26] WACV19 Olympics 1189 1.62h $\times$ - - $\times$ -
Rhythmic Gymnastics [52] ACMMM20 Olympics 1000 23.4h $\times$ - - $\times$ -
FineDiving [45] CVPR22 Olympics 3000 3.5h $\times$ - - $\times$ -
LOGO [53] CVPR23 Olympics 200 11.4h $\times$ - - $\times$ -
Skate-IRIS [22] ACMIUI23 MIT-Skate [31] 150 7.3h Score 1050 $\times$ -
CPR-Coach [39] CVPR24 Live Saving 1416 7.7h Binary 18.4k $\times$ $\times$ -
@@
MTL-AQA [27] CVPR19 Olympics 1412 1.7h $\times$ - - 1.94
MTL-NAE [54] CVPR24 MTL-AQA [27] 1412 1.7h $\times$ - - 2.94
EgoExo-Fitness [15] ECCV24 Daily Fitness 913 4.6h Binary 7.8k 2.37
EgoExo4D (Sports) [10] CVPR24 Daily Sports 1219 16.4h $\times$ - - 3.44
EE4D-DAC(Ours) EgoExo4D [10] 4843 10.8h Commentary 25.1k 4.33
Table 1: Comparison to popular Action Assessment datasets. Our EE4D-DAC dataset is the first dataset that contains detailed coaching commentary on both key-point level and instance level. Coach-Score: A metric that evaluates whether the provided commentary is suited for real-world coaching scenarios (See Sec. 3.3 for details).
@@ 3.2 Annotation Pipeline
By going deeper into the expert commentary in EgoExo4D, it can be observed that: 1) The original expert commentary is highly colloquial and noisy, awaiting summarization before being used to build a coaching model. 2) The expert commentary is highly associated with the general technical keypoints, making it possible to mine the relationships between the general technical keypoints and final coaching commentary.
Based on the observations, we propose an LLM-based automatic annotation pipeline to obtain general technical keypoints, and hierarchical coaching commentary on both keypoint level and instance level, which is shown in Fig. 2. We will bring more details as follows.
- Collection of the General Technical Keypoints. In the general coaching-learning scenario, a coach will provide several technical keypoints in advance so that the skill learner can learn to follow. Considering that completing a skilled physical action always requires cooperation across multiple body parts and objects, we define a technical keypoint as a language instruction on one of the seven dimensions, including six body parts (e.g., head & eyes, arms & hands) and the human-objects interaction.
To obtain technical keypoints for each skilled action task, we treat the LLM (GPT-4O) as a general coach. As shown in the lower left part of Fig. 2, by providing the task name (e.g., Basketball Drills - Reverse Layup) and detailed instructions to the LLM, we are able to obtain the general technical keypoints on the previously mentioned seven dimensions (e.g., For the dimension of “head & eyes”, we have the keypoint: “The player should keep the head up and eyes focused on the backboard and rim to accurately judge the distance and angle for the reverse layup.”).
- Collection of Hierarchical Coaching Commentary. Subsequently, we further ask the LLM to mine the relationships between the technical keypoints and origin expert commentary and provide summarized hierarchical (i.e., keypoint level and instance level) coaching commentary about what is done well and what can be improved in the action execution.
As shown in the right part of Fig. 2, in this phase, LLM is asked to finish two tasks: (1) Review the general technical keypoints and the original expert commentary, and then summarize the strength and weakness in the execution corresponding to each keypoint. (2) Review the original expert commentary and the generated strengths and weaknesses, and then provide the overall commentary on the execution.
To sum up, with the annotation pipeline, for each unique instance, we construct rich language annotations including the General Technical Keypoints, Keypoint-level Commentary, and Instance-level Commentary. Moreover, for each instance we also obtain an average rating by averaging all the ratings provided by different experts.
- Discussions. Note that in some cases, an execution could reveal both the strength and weakness aspects on one keypoint, and some may just reveal one of them. Such a characteristic is inherited from EgoExo4D [10].

### v2

# TechCoach: Towards Technical-Point-Aware Descriptive Action Coaching

@@ Abstract
To guide a learner in mastering action skills, it is crucial for a coach to 1) reason through the learner’s action execution and technical points (TechPoints), and 2) provide detailed, comprehensible feedback on what is done well and what can be improved. However, existing score-based action assessment methods are still far from reaching this practical scenario. To bridge this gap, we investigate a new task termed Descriptive Action Coaching (DescCoach) which requires the model to provide detailed commentary on what is done well and what can be improved beyond a simple quality score for action execution. To this end, we first build a new dataset named EE4D-DescCoach. Through an automatic annotation pipeline, our dataset goes beyond the existing action assessment datasets by providing detailed TechPoint-level commentary. Furthermore, we propose TechCoach, a new framework that explicitly incorporates TechPoint-level reasoning into the DescCoach process. The central to our method lies in the Context-aware TechPoint Reasoner, which enables TechCoach to learn TechPoint-related quality representation by querying visual context under the supervision of TechPoint-level coaching commentary. By leveraging the visual context and the TechPoint-related quality representation, a unified TechPoint-aware Action Assessor is then employed to provide the overall coaching commentary together with the quality score. Combining all of these, we establish a new benchmark for DescCoach and evaluate the effectiveness of our method through extensive experiments. The data and code will be made publicly available.
footnotetext: $\dagger$: Corresponding author. *: Equal contributions.
@@ 1 Introduction
Figure 1: An illustration of real-world action coaching. Upper: Given an action demonstration, a human coach will provide coaching comments on what is done well and what can be improved. Lower: To provide precise feedback, the coach incorporates TechPoint-level Reasoning into the action coaching process, i.e., keep the technical points (e.g., The player should extend the arms fully as they bring the ball under the basket) in mind and determine the technical-point-level action quality (e.g., The athlete performs beautifully with the right arm extension) of the action. Such a reasoning process motivates us to propose our TechCoach to address Descriptive Action Coaching.
@@
- Functionality. In practical action-guiding scenarios, a coach must provide detailed and understandable feedback on what is done well and what can be improved so that the learner can be fully aware of the execution details and master the skills (see the upper part of Fig. 1). However, existing works formulate AQA as a score-regression [31, 38, 54] or pairwise-ranking [9, 60, 13] problem, limiting the applicability of the models to more practical scenarios.
@@
- Reasoning process. To provide precise feedback, a coach will keep the technical points (TechPoints) in mind (e.g., The player should extend the arms fully in Reverse Layup) and determine the TechPoint-level action quality from action execution (see the lower part of Fig. 1). However, few existing works have incorporated such a TechPoint-level reasoning process into AQA. While recent works [42, 16] propose to address this issue by recognizing the existence of TechPoint-level mistakes, such a paradigm still limits deeper exploration of the connections between the action execution and the TechPoint-level action quality (i.e., detailed strength/weakness on a TechPoint).
@@
To bridge the gap between the field of AQA and real-world action coaching scenarios, we investigate a new task termed Descriptive Action Coaching (DescCoach), which requires the model to provide detailed commentary on what is done well and what can be improved, beyond merely assigning a quality score to an action execution. This task is challenging because to provide such a detailed commentary, the model not only needs to understand fine-grained action execution (e.g., the player looks down when performing Reverse Layup), but also establish explicit connections between the execution with the action quality (e.g., it is a weakness to look down at the ball as it will miss the target).
@@
To facilitate the studies on DescCoach, we construct EE4D-DescCoach, a new dataset that not only contains various action videos and quality scores but also features hierarchical detailed coaching commentary on both TechPoint and instance levels. Specifically, we first source action videos from the recently proposed EgoExo4D [10] dataset. Subsequently, we design an LLM-driven automatic annotation pipeline to progressively obtain the general TechPoints and hierarchical coaching commentary. To our knowledge, no existing AQA dataset provides detailed TechPoint-level coaching commentary. Compared to existing datasets with score-based or binary TechPoint-level annotations [16, 25], we believe the TechPoint-level commentary provides more clear and detailed clues for reasoning about action quality and providing instance-level coaching commentary.
@@
The construction of the EE4D-DescCoach dataset allows us to build TechPoint-Aware Descriptive Action Coach (TechCoach), a new DescCoach framework which explicitly incorporates TechPoint-level reasoning into the action coaching process under the supervision of TechPoint-level coaching commentary, and delivers feedback by incorporating the reasoning results and visual context.
@@
To this end, TechCoach includes a Context-aware TechPoint Reasoner, which queries the visual context with general TechPoints to obtain TechPoint-related quality representations. To ensure these representations carry the action quality information, a TechPoint-level alignment loss is proposed to align the representations with the TechPoint-level coaching commentary. After that, a Unified TechPoint-aware Action Assessor (TA2) is employed to regress the action quality score and generate overall coaching commentary. Throughout this process, a progressive action coaching attention mask guides TA2 to integrate information starting with the visual context level, followed by the TechPoint level, and ultimately the decision level (i.e., score-prediction and commentary-generation).
@@
Combining the EE4D-DescCoach dataset and TechCoach, we establish a new benchmark to expand the field of AQA. Extensive experiments demonstrate that: (1) Our TechCoach achieves state-of-the-art performance among the compared methods, including task-specific models and general MLLMs; (2) The Context-aware TechPoint Reasoner is a core design to enhance action quality perceiving ability and cannot be simply replaced by naive alternative solutions (e.g., replacing the TechPoint-level alignment by classifying whether the execution shows strength or weakness on each TechPoint).
@@
The contributions of this work can be summarized as: (1) We introduce EE4D-DescCoach, a new dataset specifically designed for Descriptive Action Coaching. The hierarchical coaching commentary, particularly at the TechPoint level, establishes EE4D-DescCoach as a unique resource among existing AQA datasets. (2) We develop TechCoach, a new method that effectively integrates TechPoint-level reasoning into the action coaching process through supervision from TechPoint-level coaching commentary. (3) We construct a new benchmark on DescCoach. Experiments not only validate the effectiveness of the proposed approach, but also lay the foundation for future research in this field.
@@
Figure 2: An overview of the LLM-driven automatic annotation pipeline for the EE4D-DescCoach dataset. We first prompt the LLM to generate general TechPoints for the given action task (Orange $\rightarrow$). Subsequently, we ask the LLM to summarize the TechPoint-level and instance-level commentary according to the general TechPoints and the original expert commentary (Blue $\rightarrow$). Information about various TechPoints is highlighted in colors. Zoom in for best view.
@@ 2 Related Works
Action Quality Assessment (AQA) [41, 20, 53, 62] aims at evaluating how well an action is performed. Most existing works pose AQA as a score regression task, and approach this task in many ways, e.g., introducing new formulations [46, 38, 54], augmenting human-body representations [27, 50], modeling with insufficient or continual data [55, 17, 63], exploring temporal structures[48, 3, 47], collaborating multi-modal information [32, 56, 45]. Although impressive progress has been achieved, regressing the final score is still far from real-world action coaching scenarios. MTL-AQA[30] and NAE[59] propose to generate commentary in addition to the score, but the commentary comes from live comment, focusing on action and score rather than skill guiding and coaching. NS-AQA[26] combines multiple models together with judging rules to provide formatted feedback for diving judging, but requires complex manual design on a single task. Differently, our work focuses on generating free-from commentary on what is done well and what can be improved, which is more aligned with practical coaching scenarios. To this end, we construct a new dataset that features hierarchical (especially the TechPoint-level) detailed coaching commentary. Comparisons with existing datasets are shown in Sec. 3.1. Furthermore, we propose a general framework that collaborates TechPoint-level reasoning into the coaching process. Different from existing works [25, 24, 33, 16, 42] that use binary or score-based supervised pipeline to introduce score rubric or detect body-part level mistakes, TechCoach represents the first approach to learn action quality-related representation under the supervision of TechPoint-level commentary. Experiments further validate the effectiveness of the unique design against other naive solutions.
@@
Video Captioning [37, 49, 19, 43, 22, 52] is a long-standing task that requires a model to generate a language caption with video frames as input. Early studies focus on designing stronger architectures [19, 43, 1, 36, 51] or utilizing video-language pretraining [37, 22, 52, 2, 35] to model relationships between video and language. Recently, Multimodal Large Language Models (MLLMs) [18, 44, 15, 23, 14, 7] show strong generalization ability across various video-language tasks such as video content description and video question answering. However, rare exploration has been conducted on whether these models could provide coaching commentary about what is done well and what can be improved. In this work, we provide a new testbed for today’s video-language models and reveal that they still struggle to address such a practical scenario. Besides, we propose a new TechPoint-guided coaching framework and evaluate its effectiveness through extensive experiments.
3 EE4D-DescCoach Dataset
@@
To enable studies on Descriptive Action Coaching, we establish a new dataset termed EE4D-DescCoach. EE4D-DescCoach not only contains various action videos and quality scores but also features hierarchical detailed coaching commentary on both technical-point (TechPoint) level and instance level. We will introduce the data source in Sec. 3.1, the annotation pipeline in Sec. 3.2, and provide comparisons with existing datasets in Sec. 3.3.
@@ 3.1 Data Source
We source data from the recently proposed EgoExo4D [10] dataset. In addition to ego-exo action videos, EgoExo4D also contains execution rating (1 to 10) and time-aligned free-form spoken commentary (e.g. “Yeah, see how she’s having to look down. You can’t see the basket here.") by domain experts. Such characteristics make the EgoExo4D dataset a great starting point for studying action coaching.
@@
{NiceTabular}
l|c|ccc|ccc|ccc Datasets Public. Data Unique Unique TechPoint Level Judgment   Instance Level Judgment
@@
MIT-Skate [34] ECCV14 Olympics 150 7.3h $\times$ - - ✓ $\times$ -
AQA-7 [29] WACV19 Olympics 1189 1.62h $\times$ - - ✓ $\times$ -
Rhythmic Gymnastics [57] ACMMM20 Olympics 1000 23.4h $\times$ - - ✓ $\times$ -
FineDiving [48] CVPR22 Olympics 3000 3.5h $\times$ - - ✓ $\times$ -
LOGO [58] CVPR23 Olympics 200 11.4h $\times$ - - ✓ $\times$ -
Skate-IRIS [25] ACMIUI23 MIT-Skate [34] 150 7.3h ✓ Score 1050 ✓ $\times$ -
CPR-Coach [42] CVPR24 Live Saving 1416 7.7h ✓ Binary 18.4k $\times$ $\times$ -
@@
MTL-AQA [30] CVPR19 Olympics 1412 1.7h $\times$ - - ✓ ✓ 1.94
MTL-NAE [59] CVPR24 MTL-AQA [30] 1412 1.7h $\times$ - - ✓ ✓ 2.94
EgoExo-Fitness [16] ECCV24 Daily Fitness 913 4.6h ✓ Binary 7.8k ✓ ✓ 2.37
EgoExo4D (Sports) [10] CVPR24 Daily Sports 1219 16.4h $\times$ - - ✓ ✓ 3.44
EE4D-DescCoach (Ours) EgoExo4D [10] 4843 10.8h Commentary 25.1k 4.33
Table 1: Comparison to popular Action Assessment datasets. Our EE4D-DescCoach dataset is the first dataset that contains hierarchical (especially TechCoach-level) detailed coaching commentary. Coach-Score: A metric that evaluates whether the provided commentary is suited for real-world coaching scenarios (See Sec. 3.3 for details).
@@ 3.2 Annotation Pipeline
By diving deeper into the expert commentary in EgoExo4D, it can be observed that: 1) The original expert commentary is highly colloquial and noisy, awaiting summarization before being used to build a coaching model. 2) The expert commentary is highly associated with the general TechPoints, making it possible to mine the relationships between the general TechPoints and final coaching commentary.
Based on the observations, we propose an LLM-based automatic annotation pipeline to obtain general TechPoints, and hierarchical coaching commentary on both TechPoint level and instance level, which is shown in Fig. 2. We will bring more details as follows.
- General TechPoints Collection. In the general coaching-learning scenario, a coach will provide several TechPoints in advance so that the skill learner can learn to follow. Considering that completing a skilled physical action always requires cooperation across multiple body parts and objects, we define a TechPoint as a language instruction on one of the seven dimensions, including six body parts (e.g., head & eyes, arms & hands) and the human-objects interaction.
To obtain TechPoints for each skilled action task, we treat the LLM (GPT-4o) as a general coach. As shown in the lower left part of Fig. 2, by providing the task name (e.g., Basketball Drills - Reverse Layup) and detailed instructions to the LLM, we are able to obtain the general TechPoints on the previously mentioned seven dimensions (e.g., For the dimension of “head & eyes”, we have the TechPoint: “The player should keep the head up and eyes focused on the backboard and rim to accurately judge the distance and angle for the reverse layup.”).
- Hierarchical Coaching Commentary Collection. Subsequently, we further ask the LLM to mine the relationships between the TechPoints and origin expert commentary and provide summarized hierarchical (i.e., TechPoint level and instance level) coaching commentary about what is done well and what can be improved in the action execution.
As shown in the right part of Fig. 2, in this phase, LLM is asked to finish two tasks: (1) Review the general TechPoints and the original expert commentary, and then summarize the strength and weakness in the execution corresponding to each TechPoint. (2) Review the original expert commentary and the generated strengths and weaknesses, and then provide the overall commentary on the execution.
To sum up, with the annotation pipeline, for each unique instance, we construct rich language annotations including the General TechPoints, TechPoint-level Commentary, and Instance-level Commentary. Moreover, for each instance we also obtain an average rating by averaging all the ratings provided by different experts.
- Discussions. Note that in some cases, an execution could reveal both the strength and weakness aspects on one TechPoint, and some may just reveal one of them. Such a characteristic is inherited from EgoExo4D [10].